OpenStudy (anonymous):

hi im having trouble finding the integral 1/u^2-16

OpenStudy (bahrom7893):

is u^2 -16 in the denominator? or is only u^2 in the denominator?

OpenStudy (anonymous):

u^2-16 is in the denominator

OpenStudy (bahrom7893):

okay

OpenStudy (bahrom7893):

workin on it

OpenStudy (anonymous):

k thank you

OpenStudy (anonymous):

substitue u for tan(x)

OpenStudy (bahrom7893):

use trig substitution..

OpenStudy (anonymous):

i cant because there is a minus sign

OpenStudy (bahrom7893):

Waiy

OpenStudy (bahrom7893):

U don't have to

OpenStudy (bahrom7893):

integration by parts

OpenStudy (anonymous):

not tangent sec

OpenStudy (bahrom7893):

Sorry meant partial fractions

OpenStudy (anonymous):

use sec(x)=u

OpenStudy (bahrom7893):

I would go this way: 1/(u^2-16) = 1/[(u-4)(u+4)]

OpenStudy (bahrom7893):

A/(u-4) + B/(u+4) = 1/(u^2-16) its easier.

OpenStudy (bahrom7893):

Au + 4A + Bu - 4B = 1

OpenStudy (anonymous):

i did that and i got -1/8lnu{u+4)+1/8ln(u-4) but its wrong

OpenStudy (bahrom7893):

Au + Bu +4A - 4B = 1 (A+B)u + (A-B)4 = 0*u + 1

OpenStudy (bahrom7893):

A+B = 0 (A-B) * 4 = 1

OpenStudy (bahrom7893):

so now let me check over my arithmetic.. I get confused when I have to type out math problems.

OpenStudy (anonymous):

k

OpenStudy (bahrom7893):

http://www.twiddla.com/499646 hey can u go there, its easier for me to write there..

OpenStudy (anonymous):

i get (1/8)ln(u-4) - (1/8)ln(u+4)